The management of a company is considering the investment in one of two projects available to the company. The returns on the two projects, A and B, are dependent on the sales outcome of the company. The Management has determined three potential sales outcomes for the company. The highest potential sales outcome for the company is outcome 1 is GH¢800,000. If this sales outcome were realized, project A would realize a return outcome of 30%; project B would realize a return of 20%. If outcome 2 were realized, the company’s sales level would be GH¢500,000. In this case, project A would yield 15%, and project B would yield 13%. The worst outcome 3 will result in a sales level of GH¢400,000, and return levels for projects A and B of 1% and 9% respectively.


1
Expert's answer
2021-04-15T07:14:44-0400
"0.3A+0.2B=800000"

"0.15A+0.13B=500000"

"0.01A+0.09B=400000"

"B=160000"

"A=2560000"